How long on average would it take an oxygen gas molecule (02) to diffuse a distance of 15.0 mm in air? Hint: See Example 12.24
Answer #1

Apply

time = x2 / (2D)

where

D = diffusion constant = 10-9 m2/s

x = distance = 15 mm = 15 *10-3 m

then, substitute

t = (15 *10-3)2 / 2 * 10-9

= 112.5 * 10-3 sec

= 3.125 * 10 -5 hours